OpenStudy (anonymous):

Solve for x: 5|2x − 2| + 8 = 18.

OpenStudy (welshfella):

you need to isolate the modulus part |2x − 2| first subtract 8 from both sides of the equation then divide both sides by 5

pooja195 (pooja195):

Add -8 to both sides. 5(|2x−2|)+8+−8=18+−8 5(|2x−2|)=10 Divide both sides by 5. 5(|2x−2|)/5=10/5 |2x−2|=2 now solve for x making 2 separate equations 2x-2=2 2x-2=-2 solve for x

OpenStudy (anonymous):

yes, that is correct

OpenStudy (anonymous):

there are two numbers that x could equal